# FREE: The Foundational Semantic Recognition for Modeling Environmental Ecosystems

### v4

@@ 3.3 Pre-training using physical simulations
The pre-trained LMs in the semantic recognition phase (Eq. 3) are not inherently trained on data specific to the target environmental ecosystems. As shown in Fig. 3 (b), the input features from samples in different seasons tend to be mixed in the latent embedding space of the original pre-trained DistilBERT model. This indicates that a standard LM, when applied directly, may fall short of capturing semantics from the text generated for our target task.
Tuning the LM towards the target domain requires sufficient observed data samples, which are often not available in real-world ecosystems. To address this, we turn to the simulated data $\{\mathbf{x}_{i,t},{\tilde{y}}_{i,t}\}$, generated by a physics-based model, and use them to pre-train the semantic recognition modeling component. Fig. 3 (a) shows that samples of different seasons can now be distinguished in the updated embedding space, which confirms that the LM model tuned with simulated data can better capture semantics in the textual descriptions $d_{i,t}$.
This simulation-based pre-training approach also helps in facilitating the training of the semantic recognition component to emulate general physical processes encoded in the physics-based model. This helps enhance the model’s generalizability as many physical processes generally hold across space and time. Upon completing this pre-training, we find that it requires only a few epochs of fine-tuning using true observations before reaching a quality model.
Figure 3: Distribution of data points randomly sampled from summer and winter in their embedding space visualized by t-SNE. (a) demonstrates a pronounced segregation between different seasons after domain-specific tuning. (b) shows an overlapping distribution of embeddings created by the original DistilBERT model.
@@ 4 Experiments
In this section, we present our datasets and provide a comprehensive assessment of the proposed methods. Our primary emphasis lies on predictive performance using sparse data, underscoring the efficacy of the proposed FREE framework and the simulation-based pre-training. Besides, we delve into additional experiments that validate our model’s capability to harness diverse input features and auxiliary observations. We further examine its performance under transfer learning scenarios and analyze its performance sensitivity by freezing the LM in the semantic recognition component during the fine-tuning phase.
4.1 Experimental Setup
We first introduce two datasets for evaluation: one for stream water temperature prediction and the other for annual crop yield prediction. Then, we introduce the baseline methods used in this paper and the implementation details.
Stream dataset
The dataset is collected from the Delaware River Basin, an ecologically diverse region and a watershed along the east coast of the United States. The datasets are pulled from the U.S. Geological Survey’s National Water Information System US Geological Survey (2016) and the Water Quality Portal Read et al. (2017), the largest standardized water quality data set for inland and coastal waterbodies Read et al. (2017). Observations at a specific latitude and longitude were matched to river segments that vary in length from 48 to 23,120 meters. These segments were defined by the national geospatial fabric used for the National Hydrologic Model as described by Regan and others (2018), and the river segments are split up to have at maximum a one-day water travel time. We match observations to river segments by snapping observations to the nearest stream segment within a tolerance of 250 meters, excluding those over 5,000 meters from a segment’s outlet. Segments with multiple observation sites were aggregated to a single daily average water temperature value.
We conduct our study on a subset of river segments of the Delaware River Basin, specifically focusing on the river segments that converge into the main stem of the Delaware River at Wilmington, DE. This large dataset is referred to as the Christina River Watershed (CRW), which contains 42 river segments. Separately, we also studied a distinct stream segment, named “single stream (SS)”, which has the largest number of water temperature observations. This allows us to test the capacity of models under data-sparse scenarios by intentionally withholding a fraction of training samples.
Our dataset spans a long period from October 31, 2006, to March 30, 2020 (4900 days). On a daily scale, we incorporated basic meteorological features (the day of the year, rainfall, daily average air temperature, and solar radiation). In addition to these, we also considered other features, namely average cloud cover fraction, groundwater temperature, subsurface temperature, and potential evapotranspiration.
While daily average water temperature observations were present for certain segments, they were not consistently available for every date. The count of temperature readings for each segment varied, ranging from a minimum of 0 to a maximum of 4,682. In totality, “CRW” comprises 205,800 temperature observations distributed across all its segments and dates, and “SS” has 4,676 temperature observations. Besides true observations, we also use the simulated water temperature data for CRW and SS generated by the physics-based PRMS-SNTemp model Markstrom (2012). For the following experiments, the pre-training process used simulated data from October 31, 2006, to March 30, 2020. We train each ML model using data from the first 2,450 days (October 31, 2006 to July 8, 2013) and then test in the next 2,450 days (July 8, 2013 to March 30, 2020).
Stream baselines
We compare model performance to multiple state-of-the-art baselines developed for stream modeling, including the Basic LSTM model (LSTM) (considered to be the state of the art in many environmental modeling problems Shen and Lawson (2021)), TransformerVaswani et al. (2017), Recurrent Graph Model (RGRN) Jia et al. (2021b), Graph-temporal convolutional network (Gr-CNN) Sun et al. (2021), HydroNets Moshe et al. (2020). Besides, we include a baseline method that assimilates new observations for updating the model, the Heterogeneous Recurrent Graph Networks with data assimilation (HRGN-DA) Chen et al. (2021a).
Table 1: Prediction RMSE for stream water temperature using 1%, 2%, 4%, and 100% randomly selected training labels. The best results are bold, while the second best results are underlined.
@@ 4.1 Experimental Setup
Crop dataset
We use the corn yield data in Illinois and Iowa from the years 2000-2020 provided by USDA National Agricultural Statistics Service (NASS)  [^1] https://quickstats.nass.usda.gov/. In particular, there are in total 199 counties in our study region (100 counties in Illinois and 99 counties in Iowa). The corn yield data (in gCm-2) are available for each county each year. The observed annual crop yield labels cover 199 countries spanning 21 years from 2000-2020, with daily records for each year. The input data include seven features, namely, the surface downward shortwave radiation, the maximum air temperature, the maximum humidity, the wind speed, the precipitation, the depth-weighted averaged bulk density in the soil, and the depth-weighted averaged sand content in the soil. The feature data collected from NLDAS-2 climate data Xia and others (2012), 0-30cm gSSURGO soil properties [^2] https://gdg.sc.egov.usda.gov/, and the crop type information.
Besides the true observed yield labels, we use the physics-based Ecosys model Zhou et al. (2021) to simulate ecosystem autotrophic respiration (Ra), ecosystem heterotrophic respiration (Rh), net ecosystem exchange (NEE), and crop yield for 10,335 synthetic sample locations in the United States from the years 2000-2020. All synthetic data are used for pre-training. To reduce computational load, we randomly sample 1/7 of synthetic data in the pre-training phase. We use the observational data from 2000-2017 for training and the data from 2018-2020 for testing.
Crop baselines
We compare the proposed method with the LSTM-based model and two other baselines that have both proven effective in crop yield prediction, i.e., the physics-guided attention network (PG-AN) He et al. (2023), and the graph neural network-based model Fan et al. (2022) pre-trained with the same simulated data (PG-GNN).
Implementation details
We pre-train the FREE model with 800 epochs and fine-tune it with an additional 50 epochs. We implemented our methods using PyTorch == 2.0.1, on an RTX 4090 GPU and employed an ADAM optimizer with a 1e-5 learning rate for DistilBERT, and 0.00025/0.0001 pre-training/fine-tuning learning rates for LSTM. Additionally, the batch size is 1 and the hidden dimension is 256.
@@ 4.2 Overall predictive performance
Table 1 and Table 2 show the performance of various methods for predicting stream water temperature and predicting annual crop yield, respectively. The notation FREE represents a model pre-trained on synthetic data and then fine-tuned on observational data. We also conduct the sparsity test on both datasets. For this sparsity test, training was conducted using only {4%, 2%, 1%} and {50%, 20%, 10%} of training data randomly sampled from the whole training period on the stream dataset and the crop dataset, respectively.
According to Table 1, the FREE outperforms all baseline methods when data are abundant, which shows the effectiveness of the proposed method in capturing the semantics and relationships of input features from textual descriptions. Notably, the FREE maintains good performance regardless of label volume, while other models generally have much worse performance in data-sparse scenarios. This underscores the model enhancement brought by the simulation-based pre-training in learning generalizable and domain-specific semantics. Moreover, pre-training our model before fine-tuning significantly reduces the time for effective fine-tuning (about one-tenth of the time compared to direct fine-tuning). Table 2 presents similar comparisons between the proposed method and the baselines in predicting annual corn yield. We also observe that some methods can have slightly better performance with reduced training data. This is because the data samples are highly heterogeneous across different counties and years in this dataset, and the randomly sampled training data have a chance to get slightly closer to the testing data in distribution.
Figure 4: Comparison of pre-trained FREE models and the baseline LSTM on stream water temperature prediction. (a) SS test case with full training data. (b) SS test case with 1% training data. (c) an example segment in CRW with full training data from CRW. (d) the same example segment in CRW with 1% training data from CRW.
Figure 5: Evaluation of FREE with auxiliary information. the variation in test errors for (a) and (b)  predicting stream water temperature upon the inclusion of extra observations under different sparsity levels in SS and CRW separately, (c) predicting stream water temperature upon the integration of additional features, (d) predicting corn yield upon the integration of additional features.
In Fig. 4, we show the predicted water temperature over one year for the SS test case and another river segment within the CRW test case to demonstrate the alignment of predictions and true observations. The figure clearly shows that our method outperforms the baseline model under different data sparsity scenarios (100% and 1%), especially in capturing the water temperature oscillations in the summertime. The LSTM model’s predictions, while reasonably tracking the general trend, fall short in replicating the nuanced variations, especially as the training data becomes sparser. In scenarios with limited data, the LSTM struggles to recognize subtle changes and creates smooth predictions, overlooking smaller peaks and valleys that our method consistently identifies and traces.
@@ 4.3 Handling auxiliary information
We also evaluate the performance of our method in harnessing auxiliary observations and diverse input features. FREE-C refers to the model that incorporates auxiliary observed water temperatures of the current river segment from the prior day (when they are available). As shown in Fig. 5 (a) and (b), FREE-C outperforms FREE, demonstrating that our proposed approach can process auxiliary observation data effectively. It is worth mentioning that in Fig. 5 (b), the baseline HRGN-DA, which also incorporates new observational data, initially achieves close performance with the proposed FREE-C approach using the complete training dataset. Nonetheless, when the available labeled data is reduced, our FREE-C models surpass HRGN-DA in accuracy. Additionally, HRGN-DA is designed for application to a graph of streams, making it incompatible with the SS dataset. For crop yield prediction, unlike daily stream temperature prediction, crop yield labels are available on a yearly scale. The yield labels from the previous year usually have little impact on the yield of the next year and thus we did not include crop yield’s auxiliary observations experiments.
[…]
Referencing Fig. 5, two primary observations are evident: (1) The integration of auxiliary observations improves the performance of the stream modeling dataset. Notably, FREE-C achieves the best results overall in both the CRW and SS datasets. (2) For both datasets, the use of additional features can reduce the prediction errors. We will explore using more diverse input features in future work.
Figure 6: Comparison between FREE, the non-pre-trained FREE, and the transferred FREE (pre-trained from CRW) under different sparsity levels in the target SS. (a) train on meteorological features. (b) trained on meteorological and four additional features.
@@ 4.4 Evaluation on model transferability
We assess the performance of adapting a model pre-trained using simulated data from CRW to the river segment in a different region (SS). In particular, we compared the transferred models (FREEtrs and FREE-A4trs) to the models trained on observational data directly without pre-training (FREEnprt and FREE-A4nprt) and the models pre-trained on the same segment in SS (FREE, FREE-A4). Here we use the subscript “trs” to represent the model being pre-trained on CRW. As Fig. 6 shows, The pre-trained models (FREEtrs, FREE-A4trs, FREE, FREE-A4) consistently outperform the models FREEnprt, FREE-A4nprt under different data sparsity levels, maintaining stable performance despite reduced label availability. Remarkably, the pre-trained model from the different domain CRW can have a similar degree of contribution to performance enhancement compared to the model pre-trained on the same target segment. These findings emphasize the pre-trained models’ ability to not only address the challenges caused by sparse labels but also to transfer learning capabilities. This shows the potential of building a global pre-trained model over large regions, which can then be adapted efficiently to different target regions with small training data.
4.5 Fine-tuning analysis
We investigate the effect of two different fine-tuning strategies: (1) fine-tuning all the parameters of the semantic recognition model, i.e., LM+LSTM (FT-All), and (2) fine-tuning only the parameters of the LSTM layer while fixing the embeddings created by the LM model (FT-LSTM). The underlying assumption for the fine-tuning approach (2) is that the LM can generate semantics-aware embeddings after being pre-trained on simulations, eliminating the need to be further adjusted on true observations. The results presented in Table 3 show that comprehensive fine-tuning FT-All over all the parameters within our pre-trained models yields better results. Interestingly, the partial fine-tuning approach FT-LSTM, where we fix the parameters of the LM and only adjust those of the LSTM, also delivers decent performance. While this tuning approach (FT-LSTM) slightly compromises the predictive accuracy compared to the comprehensive fine-tuning approach (FT-All), it requires much less computational cost, making it a convenient and practical alternative in resource-limited scenarios.
Table 3: Prediction RMSE for stream water temperature and crop yield using different fine-tuned models. FT-All means fine-tuning the whole model using training labels, and FT-LSTM means only fine-tuning the LSTM layer.
Dataset Method FT-All FT-LSTM
SS FREE 1.70 1.71
FREE-C 1.11 1.70
FREE-A2 1.73 1.98
FREE-A3 1.68 1.93
FREE-A4 1.63 1.94
CRW FREE 1.61 2.22
FREE-C 1.02 2.35
FREE-U 1.50 2.28
FREE-CU 1.03 2.27
FREE-A2 1.68 2.18
FREE-A3 1.61 1.98
FREE-A4 1.65 2.11
Crop FREE-A0 65.70 66.73
FREE-A1 65.45 66.68
FREE-A2 65.24 66.01
@@ 5 Conclusion
This paper introduces FREE, a novel LLM-based framework for modeling environmental ecosystems. FREE is shown to outperform existing methods in the context of predicting stream water temperature and annual crop yield. This demonstrates the potential of FREE in better capturing the semantics of environmental features by converting them to natural language descriptions. FREE also has the flexibility to utilize different input features and auxiliary observations for different data samples to enhance predictive performance. Moreover, FREE maintains high predictive performance even under data-sparse scenarios. This is attributed to the model initialization through the pre-training on simulated data. The simulation-based pre-training also aids in extracting physically consistent data patterns, which improves the generalizability and transferability to different regions.
We anticipate the FREE framework to serve as a stepping stone to building foundational solutions for modeling complex environmental and physical systems. This framework remains limited in (1) its computational cost (e.g., caused by the increasing length of textual description) when combining a large number of physical variables as inputs, (2) the ability to integrate known physical and causal relationships in the complex ecosystems, and (3) effectively handling highly erratic or noisy data beyond incomplete or missing data scenarios. We will investigate these directions in future work.

### v5

@@ 2.3 Pre-training using physical simulations
Figure 3: t-SNE of embeddings for data points randomly sampled from summer and winter.
The LM in the semantic recognition phase (Eq. 3) are not inherently trained on data specific to the target environmental ecosystems. As shown in Fig. 3 (b), the input features samples from different seasons tend to be mixed in the latent embedding space of the original LM, when applied directly, may fall short of capturing semantics from the text generated for our target task.
Tuning the LM towards the target domain requires sufficient observed samples, which are often not available in real-world ecosystems. Therefore, we use simulated data $\tilde{y}_{i,t}$ generated by physics-based model to pre-train the semantic recognition component. Fig. 3 (a) shows that samples of different seasons can now be distinguished, confirming that the LM model tuned with simulated data can better capture semantics in the textual descriptions $d_{i,t}$.
This simulation-based pre-training also facilitates the training of the semantic recognition component to emulate general physical processes encoded in the physics-based model, enhancing the model’s generalizability as many physical processes generally hold across space and time. Upon completing this pre-training, it requires only a few epochs of fine-tuning using true observations before reaching a quality model.
@@ 3 Experiments
Table 1: Prediction RMSE for stream temperature using 1%, 2%, 4%, and 100% randomly selected training labels. The best results are bold, the second best results are underlined.
@@
In this section, we present our datasets and provide a comprehensive assessment of the proposed methods. Our primary emphasis lies on predictive performance using sparse data, underscoring the efficacy of the proposed FREE framework and the simulation-based pre-training. Besides, we delve into additional experiments that validate our model’s capability to harness diverse input features and auxiliary observations. We further examine its performance under transfer learning scenarios.
3.1 Experimental Setup
Stream dataset and baselines
We use data from the Delaware River Basin (DRB), an ecologically diverse region and a watershed along the east coast of US. We study two subsets: Christina River Watershed (CRW) (with 42 connected river segments), and a single stream segment from a distinct region (SS). We train each ML model using data from October 31, 2006 to July 8, 2013 (2,450 days), and test on the following 2,450 days, up to March 30, 2020. On a daily scale, we incorporated basic meteorological features (the day of the year, rainfall, daily average air temperature, and solar radiation). In addition to these, we also considered other features, namely average cloud cover fraction, groundwater temperature, subsurface temperature, and potential evapotranspiration. We compare FREE with multiple popular baselines developed for stream modeling, including LSTM model (considered to be the state of the art in many environmental modeling problems (Shen & Lawson, 2021)), Transformer (Vaswani et al., 2017), RGRN (Jia et al., 2021),Gr-CNN (Sun et al., 2021), HydroNets (Moshe et al., 2020). Besides, we include a baseline that assimilates new observations for updating the model, HRGN-DA (Chen et al., 2021).
Crop dataset and baselines
We use the corn yield data in Illinois and Iowa from the years 2000-2020 provided by USDA National Agricultural Statistics Service (NASS) (available at https://quickstats.nass.usda.gov/). The observed annual crop yield labels cover 199 countries spanning 21 years from 2000-2020, with daily records for each year. The input data include seven features: the surface downward shortwave radiation, air temperature, humidity, wind speed, precipitation, depth-weighted averaged bulk density in the soil, and depth-weighted averaged sand content in the soil. To reduce computational load, we randomly sample 1/7 of synthetic data in the pre-training phase. We use the observational data from 2000-2017 for training and the data from 2018-2020 for testing. We compare the proposed method with the LSTM and other two baselines that have both proven effective in crop yield prediction, i.e., PG-AN (He et al., 2023), and PG-GNN (Fan et al., 2022).
@@ 3.2 Overall predictive performance
Figure 4: Comparison of FREE and LSTM on stream water temperature prediction.
Tables 1 and 2 show the performance of various methods for predicting stream water temperature and annual crop yield, respectively. The notation FREE represents a model pre-trained on synthetic data and then fine-tuned on observational data. Table 1 shows that FREE outperforms all baselines when data are abundant, showing its effectiveness in capturing the semantics and relationships of input features from textual descriptions. Notably, the FREE maintains good performance regardless of label volume, while other models generally have much worse performance in data-sparse scenarios. This underscores the model enhancement brought by the simulation-based pre-training in learning generalizable and domain-specific semantics. Moreover, pre-training our model significantly reduces the time for effective fine-tuning (about one-tenth of the time compared to direct fine-tuning). Fig. 4 shows the predicted water temperature over one year for the SS and a river segment within the CRW to demonstrate the alignment of predictions and true observations. The figure clearly shows that our method outperforms the baseline model under different data sparsity scenarios (100% and 1%), especially in capturing the water temperature oscillations in the summertime. The LSTM’s predictions, while reasonably tracking the general trend, fall short in replicating the nuanced variations, especially as the training data becomes sparser. LSTM struggles to recognize subtle changes and creates smooth predictions with limited data, overlooking smaller peaks and valleys that our method consistently identifies and traces.
@@ 3.3 Handling auxiliary information
We evaluate FREE’s ability to harness auxiliary observations and diverse input features. FREE-C refers to the model that incorporates auxiliary observed water temperatures of the current river segment from the prior day (when they are available). We show in Fig. 5 (a) and (b) the test errors for predicting stream water temperature upon the inclusion of extra observations under different sparsity levels in SS and CRW, respectively. FREE-C outperforms FREE, demonstrating that our proposed approach can process auxiliary observation data effectively. It is worth mentioning that in Fig. 5 (b), the baseline HRGN-DA (it is designed for application to a graph of streams, making it incompatible with the SS dataset), which also incorporates new observational data, initially achieves a close performance with FREE-C using the complete training dataset. Nonetheless, when the available labeled data is reduced, FREE-C surpasses HRGN-DA in accuracy. Unlike daily stream temperature prediction, crop yield labels are available on a yearly scale. The yield labels from the previous year usually have little impact on the yield of the next year, thus we did not include crop yield experiments here. On the other hand, we consider variants of FREE using different sets of features. To mimic different input features in real scenarios, we create data samples with the following assumption.
Figure 5: Evaluation of FREE with auxiliary information.
All data samples have access to the meteorological features as the daily weather data are publicly available. Different data samples may have different features from other data sources, which are randomly sampled. We use FREE-Amm to represent the variant of the FREE method that uses the meteorological features and $m$ randomly selected additional features. Fig. 5 (c) shows the performance of predicting water temperature upon the integration of additional features, and Fig. 5 (d) shows the predictive performance of corn yield upon the integration of additional features. For both datasets, the use of additional features can reduce the prediction errors.
@@ 3.4 Evaluation on model transferability
Figure 6: Comparison of RMSE under different sparsity levels in the target $S$. (a) train on meteorological features. (b) trained on meteorological and four additional features.
We assess transfer learning by adapting CRW-pretrained models to the SS region. Specifically, we compare the transferred models pre-trained on CRW (FREEtrs and FREE-A4trs) with models pre-trained on the same segment in SS (FREE, FREE-A4), and with those trained on observational data directly without pre-training (FREEnprt and FREE-A4nprt). As Fig. 6 shows, the pre-trained models consistently outperform the models without pre-training under different data sparsity levels, maintaining stable performance despite reduced label availability. Remarkably, the pre-trained model from a different domain can have a similar degree of contribution to performance enhancement compared to the model pre-trained on the same target segment. These findings emphasize the potential of this approach for building a global pre-trained model over large regions with small training data.
@@ 4 Conclusion
This paper introduces FREE, a novel LLM-based framework for environmental modeling that converts different input features and auxiliary observations for different data samples into natural language for semantic recognition. FREE is shown to outperform existing methods in the context of predicting stream water temperature and annual crop yield, especially under data sparsity. The simulation-based pre-training also aids in extracting physically consistent data patterns, which improves the generalizability and transferability to different regions. We anticipate the FREE framework to serve as a stepping stone to building foundational solutions for modeling complex environmental and physical systems.
5 Acknowledgments
This work was supported by the National Science Foundation (NSF) under grants 2239175, 2316305, 2147195, 2203581, 2425844, 2425845, 2430978, 2126474, 2530609, and 2530610; the USGS awards G21AC10564 and G22AC00266; the NASA grants 80NSSC24K1061 and 80NSSC25K0013; and the NSF NCAR’s Derecho HPC system. This research was also supported in part by the University of Pittsburgh Center for Research Computing through the resources provided.